The coffee counter charges $11.00 per pound for Kenyan coffee and $13.00 per pound for Sumatran coffee.
How much of each type should be used to make a 20 pound blend that sells for $12.00 per pound?
The coffee counter should mix _____pounds of Kenyan coffee and ______pounds of Sumatran coffee to make 20 pounds of a blend that sells for $12.00 per pound.

Answer by richwmiller(17219)

11k+13s=20*12,
k+s=20
k = 10, s = 10